OpenStudy (anonymous):

a+8=4 _______ -2.1 solve then check solution

OpenStudy (anonymous):

help

jimthompson5910 (jim_thompson5910):

The equation is \[\Large \frac{a+8}{-2.1} = 4\] right???

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

\[\Large \frac{a+8}{-2.1} = 4\] \[\Large a+8 = 4(-2.1)\] \[\Large a+8 = -8.4\] Does that help?

jimthompson5910 (jim_thompson5910):

I multiplied both sides by -2.1

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

Tell me what you get for the answer